# Amortized SHAP values via sparse Fourier function approximation

### v1

@@ Abstract
SHAP values – a.k.a. SHapley Additive exPlanations – are a popular local feature-attribution method widely used in interpretable and explainable AI. We tackle the problem of efficiently computing these values. We cover both the model-agnostic (black-box) setting, where one only has query access to the model and also the case of (ensembles of) trees where one has access to the structure of the tree. For both the black-box and the tree setting we propose a two-stage approach for estimating SHAP values.
Our algorithm’s first step harnesses recent results showing that many real-world predictors have a spectral bias that allows us to either exactly represent (in the case of ensembles of decision trees), or efficiently approximate them (in the case of neural networks) using a compact Fourier representation. For the case of trees, given access to the tree structure, one can extract the Fourier representation using a simple recursive algorithm. For the black-box setting, given query access to the black-box function, we utilize a sparse Fourier approximation algorithm to efficiently extract its compact Fourier approximation.
In the second step of the algorithm, we use the Fourier representation to exactly compute SHAP values. The second step is computationally very cheap because firstly, the representation is compact and secondly, we prove that there exists a closed-form expression for SHAP values for the Fourier basis functions. Furthermore, the expression we derive effectively “linearizes” the computation into a simple summation and is amenable to parallelization on multiple cores or a GPU. Since the function approximation (first step) is only done once, it allows us to produce Shapley values in an amortized way. We show speedups compared to relevant baseline methods equal levels of accuracy for both the tree and black-box settings. Moreover, this approach introduces a reliable and fine-grained continuous trade-off between computation and accuracy through the sparsity of the Fourier approximation, a feature previously unavailable in all black-box methods.
@@ 1 Introduction
Interpretability of machine learning models is paramount, especially in high-stakes applications in areas such as medicine, fraud detection, or credit scoring. This is crucial to the extent that in Europe, the General Data Protection Regulation (GDPR) mandates the legal right to an explanation of algorithmic decisions (Voigt & Von dem Bussche, 2017). Say we are given a predictor/model $h:\mathcal{X}^{n}\rightarrow\mathbb{R}$ which maps an input (data) instance $x^{*}\in\mathcal{X}^{n}$ to a prediction $h(x^{*})$. Instance-wise a.k.a. local feature attribution methods assign “importances” to each of the features $x_{i}^{*}\in\mathcal{X}$ of the instance $x^{*}$ which quantify how influential that feature was in the model predicting the value $h(x^{*})$.
@@
A widely used method for deriving attributions (importances) is the notion of SHapley Additive exPlanations, commonly referred to simply as SHAP values. Originally, the notion of Shapley values was introduced in the seminal work of Shapley (1952) in the context of cooperative game theory. The Shapley value is a mathematically well-founded and “fair” way of distributing a reward among all the members of a group playing a cooperative game and it is computed based on the rewards that would be received for all possible coalitions. The Shapley value is the unique way of distributing the reward that satisfies several reasonable mathematical properties that capture a notion of fairness (Shapley, 1952). In the context of machine learning and statistics, the players become features, the reward is the prediction of the predictor $h$ and the SHAP value is the “contribution” or “influence” of that feature on the prediction. Shapley values are widely used due to their mathematical soundness and desirable properties Gromping (2007); Štrumbelj et al. (2009); Owen (2014); Datta et al. (2016); Owen & Prieur (2017); Lundberg & Lee (2017); Lundberg et al. (2020) Aas et al. (2021).
@@
Despite their prevalence, the computation of SHAP values is challenging, since it involves an exponential sum (considering the importance of a feature in the context of all possible “coalitions” of other features). Therefore, approximating them and speeding up the computation has received attention in a variety of settings. SHAP value computation can easily dominate the computation time of industry-level machine learning solutions on datasets with millions or more entries (Yang, 2021). Yang (2021) point out that industrial applications sometimes require hundreds of millions of samples to be explained. Examples include feed ranking, ads targeting, and subscription propensity models. In these modeling pipelines, spending tens of hours in model interpretation becomes a significant bottleneck (Yang, 2021) and one usually needs to resort to multiple cores and parallel computing.
@@
Significant work has gone into speeding up the computation of SHAP values for a variety of settings. In the (ensemble of) trees setting, full access to the tree structure is assumed. Yang (2021); Bifet et al. (2022) provide theoretical and practical computational speedups to the classic TreeShap (Lundberg et al., 2020). Similar to these results, in this work, we provide significant speedups for the tree setting over previous methods.
@@
As opposed to the tree setting, which is a “white box” setting, in the model-agnostic a.k.a black-box setting, we only have query access to the model. Here, our only means of access to the predictor is that we can pick an arbitrary $x\in\mathcal{X}^{n}$ and query the predictor for its value $h(x)$. The usual approach here is to approximate the exponential sum of the SHAP value computations using stochastic sampling (Covert & Lee, 2020; Mitchell et al., 2022a; Lundberg & Lee, 2017). In this setting, Covert & Lee (2020); Mitchell et al. (2022a) provide sampling methods that require fewer queries to the black-box compared to vanilla KernelShap (Lundberg & Lee, 2017) for equal approximation accuracy. FastShap Jethani et al. (2021), which introduces a method for estimating Shapley values in a single forward pass using an end-to-end learned explainer neural network model. Our algorithm FourierShap falls into the query-access black-box setting as well.
@@
However, we take a different approach. We are guided by the key insight that many models used in practice have a “spectral bias”. Yang & Salman (2019); Valle-Perez et al. (2018) provably and experimentally show that fully connected neural networks with binary (zero-one) inputs learn low-degree – and therefore sparse – functions in a basis called the Walsh-Hadamard a.k.a Fourier basis. It is well known that the Walsh-Hadamard transform (WHT) of an ensemble of $T$ trees of depth $d$ is also of degree at most $d$ and moreover, $k=O(T4^{d})$-sparse (Kushilevitz & Mansour, 1993; Mansour, 1994).
@@
Our contributions: Guided by the aforementioned insights, we provide an algorithm to compute SHAP values, using the Fourier representation of the tree or black-box model. We first approximate the black-box function by taking its sparse Fourier transform. We theoretically justify, and show through extensive experiments, that for many real-world models such as fully connected neural networks and (ensembles of) trees this representation is accurate. Subsequently, we prove that SHAP values for a single Fourier basis function admit a closed-form expression not involving an exponential summation. Therefore, using the Fourier representation we overcome the exponential sum and can utilize compute power effectively to compute SHAP values. Furthermore, the closed-form expression we derive effectively “linearizes” the computation into a simple summation and is amenable to parallelization on multiple cores or a GPU. The Fourier approximation step is only done once, therefore FourierShap amortizes the cost of computing explanations for many inputs. Subsequently, SHAP computations using the Fourier approximation are orders of magnitude faster compared to other black-box SHAP approximation methods such as KernelShap and other variations of it which all involve a computationally expensive optimization. We show speedups over other methods such as DeepLift (Shrikumar et al., 2017) and FastShap (Jethani et al., 2021) as well. In addition to the speedup, our algorithm enables a reliable continuous trade-off between computation and accuracy, controlled in a fine manner by the sparsity of the Fourier approximation. This was not formerly possible with these two black box methods.
@@ 2.1 Shapley values
In game theory, a cooperative game is a function $v:2^{[n]}\rightarrow\mathbb{R}$ that maps a subset (coalition) $S\subseteq[n]$ of a group of players $[n]=\{1,\dots,n\}$ to their total reward (when they cooperate). If all the players cooperate they win a total reward of value $v([n])$. The main question is how they would distribute this reward among themselves. Shapley (1952) resolved this problem by a deriving a unique value based on “fairness axioms” proposed in his seminal work (Shapley, 1952). The Shapley value of player $i\in[n]$ is:
@@
In the machine learning context, we have a predictor $h:\mathcal{X}^{n}\rightarrow\mathbb{R}$ mapping an $n$-dimensional feature vector to a value. In this context, the players become features $x_{i}\in\mathcal{X}$ and the reward is the prediction of the predictor $h$ and the Shapley value is the “contribution” or “influence” of the $i$’th feature on the prediction. We define $v(S)$ accordingly to capture this notion (Lundberg & Lee, 2017):
@@
The choice of which distribution to average the missing features from has been discussed thoroughly in the relevant literature. As mentioned before, in this work, we focus on the SHAP values as defined in KernelShap introduced by Lundberg & Lee (2017); Lundberg et al. (2020) also know as “Interventional” (Janzing et al., 2020; Van den Broeck et al., 2021) or “Baseline” (Sundararajan & Najmi, 2020) SHAP, where the missing features are integrated out from the marginal distribution $p(\bm{x_{[n]\setminus S}})$, as opposed to the conditional distribution $p(\bm{x_{[n]\setminus S}}|\bm{x_{S}})$. We refer the reader to Appendix B.3 for a comprehensive overview of the literature discussing these two notions and their conceptual differences.
Since we will be using the well-known KernelShap (Lundberg & Lee, 2017) and its variant LinRegShap (Covert & Lee, 2020) as a baseline we briefly review their method here. Lundberg & Lee (2017) propose the “least squares characterization” of SHAP values. They prove that SHAP values are the solution to the following minimization problem:
@@
Unfortunately, the above optimization still involves an exponential sum. Therefore, Lundberg & Lee (2017) propose to sample subsets $S$ uniformly at random. Covert & Lee (2020); Mitchell et al. (2022a) provide better ways of sampling and solving the optimization to get approximations with lower variances and biases. Nevertheless, all these methods require solving a least squares minimization subject to constraints for each explained instance $x^{*}$ and, therefore, are computationally expensive.
Later, we also compare our method to FastSHAP (Jethani et al., 2021), a model-agnostic algorithm for computing SHAP values. In FastSHAP, a parametric explainer $\phi$ (e.g., MLP) is trained to directly generate SHAP values given data samples. Since training is only done once, this, similar to us, amortizes the cost of generating SHAP values across multiple instances. Computing SHAP values only requires a forward pass on the trained model and therefore can be done very quickly.
@@ 2.2 Fourier representations
We will later use a sparse Fourier representation to approximate the black-box function. Here we review the notions of the Fourier basis and what we mean by sparsity. In this work, we focus on the setting where the inputs to the black-box function (predictor) are binary, i.e., $\mathcal{X}^{n}=\{0,1\}^{n}$. This means, we assume we have binary features and/or categorical features, through standard one-hot representations  [^1] Continuous features can be discretized into quantiles, enabling their transformation into categorical features.. The Fourier representation of the pseudo-boolean function $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ is the unique expansion of $h$ as follows: $h(x)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{f\in\{0,1\}^{n}}\widehat{h}(f)(-1)^{\langle f,x\rangle}$.
The inner product of two vectors $f,x\in\{0,1\}^{n}$ is defined as: $\langle f,x\rangle\equiv\sum_{i}^{n}f_{i}x_{i},\forall x,y\in\{0,1\}^{n}$.
@@
Assume $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ can be decomposed as follows: $h(x)=\sum\limits_{i=1}^{p}h_{i}(x_{S_{i}}),S_{i}\subseteq[n]$. That is, each function $h_{i}:\{0,1\}^{|S_{i}|}\rightarrow\mathbb{R}$ depends on at most $|S_{i}|$ variables. Then, $h$ is $k=O(\sum\limits_{i=1}^{p}2^{|S_{i}|})$-sparse and of degree $d=\max(|S_{1}|,\dots,|S_{p}|)$. (Proof in Appendix C.1)
@@
Let, $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be a pseudo-boolean function and let $d\in\mathbb{N}$ be some constant (w.r.t. $n$). If $h$ is of degree $d$, then, it is $k=O(n^{d})$-sparse. (Proof in Appendix C.1)
@@ 3.1 Spectral bias of fully connected neural networks
The function a fully connected neural network represents at initialization is a sample from a Gaussian Process (GP) (Rasmussen, 2004) in the infinite-width limit. Here, the randomness is over the initial weights and biases. The kernel $K$ of this GP is called the Conjugate Kernel (CK) (Daniely et al., 2016; Lee et al., 2017). A mentioned before, here we investigate the case where the inputs to the neural network are binary, i.e., $\mathcal{X}^{d}=\{0,1\}^{d}$ similar to (Yang & Salman, 2019; Valle-Perez et al., 2018). The CK kernel Gram matrix formed on the whole input space $\mathcal{X}^{d}=\{0,1\}^{d}$, has a simple eigenvalue decomposition :
$\mathcal{K}\in\mathbb{R}^{2^{d}\times 2^{d}},\mathcal{K}=\sum\limits_{f\in\{0,1\}^{n}}\lambda_{f}u_{f}u_{f}^{\top}\;\;\;\;$
@@
where $u_{f}\in\mathbb{R}^{2^{d}}$ is the eigenvector formed by evaluating the Fourier basis function $\Psi_{f}$ for different values of $x\in\{0,1\}^{n}$. Moreover, Yang & Salman (2019) show a weak spectral bias result in terms of the degree of $f$. Namely, the eigenvalues corresponding to higher degree frequencies have smaller values  [^2] To be more precise, they show that the eigenvalues corresponding to even and odd degree frequencies form decreasing sequences. That is, even and odd degrees are considered separately.. Given the kernel, $K$, and viewing a randomly initialized neural network function evaluated on the boolean $\{0,1\}^{n}$ as a sample from a GP one can see the following: This sample, roughly speaking, looks like a linear combination of the eigenvectors with the largest eigenvalues. This is due to the fact that a sample of the GP can be obtained as $\sum\limits_{i=1}^{2^{n}}\lambda_{i}\bm{w_{i}}u_{i},\bm{w_{i}}\sim\mathcal{N}(0,1)$. Combining this with the spectral bias results implies that neural networks are low-degree functions when randomly initialized.
@@
Going beyond neural networks at initialization, numerous studies have investigated the behavior of fully connected neural networks trained through (stochastic) gradient descent. Chizat et al. (2019); Jacot et al. (2018b); Du et al. (2018); Allen-Zhu et al. (2019a; b) found that the weights of infinite-width neural networks after training do not deviate significantly from their initialization, which has been dubbed ”lazy training” by Chizat et al. (2019). Lee et al. (2018; 2019) showed that training the last layer of an infinite-width randomly initialized neural network for an infinite amount of time corresponds to Gaussian process (GP) posterior inference with a certain kernel. Jacot et al. (2018b) extended the aforementioned results to training all the layers of a neural network (not just the final layer). They showed the evolution of an infinite-width neural network function can be described by the “Neural Tangent Kernel” (NTK, Jacot et al., 2018a) with the trained network yielding, on average, the posterior mean prediction of the corresponding GP after an infinite amount of training time. Lee et al. (2019) empirically showed the results carry over to the finite-width setting through extensive experiments. Yang & Salman (2019) again showed that the $u_{f}\in\mathbb{R}^{2^{d}}$ defined above are eigenvectors of the NTK Gram matrix and spectral bias holds. Gorji et al. (2023) validated these theoretical findings through extensive experiments in finite-width neural networks by showing that neural networks have “less tendency” to learn high-degree frequencies. We refer the reader to Appendix B.1 for a more comprehensive review.

### v2

# SHAP values via sparse Fourier representation

@@ Abstract
SHAP (SHapley Additive exPlanations) values are a widely used method for local feature attribution in interpretable and explainable AI. We propose an efficient two-stage algorithm for computing SHAP values in both black-box setting and tree-based models. Motivated by spectral bias in real-world predictors, we first approximate models using compact Fourier representations, exactly for trees and approximately for black-box models. In the second stage, we introduce a closed-form formula for exactly computing SHAP values using the Fourier representation, that “linearizes” the computation into a simple summation and is amenable to parallelization. As the Fourier approximation is computed only once, our method enables amortized SHAP value computation, achieving significant speedups over existing methods and a tunable trade-off between efficiency and precision.
@@ 1 Introduction
Interpretability of machine learning models is paramount, especially in high-stakes applications in areas such as medicine, fraud detection, or credit scoring. This is crucial to the extent that in Europe, the General Data Protection Regulation (GDPR) mandates the legal right to an explanation of algorithmic decisions [Voigt and Von dem Bussche, 2017]. Say we are given a predictor/model $h:\mathcal{X}^{n}\rightarrow\mathbb{R}$ which maps an input (data) instance $x^{*}\in\mathcal{X}^{n}$ to a prediction $h(x^{*})$. Instance-wise a.k.a. local feature attribution methods assign “importances” to each of the features $x_{i}^{*}\in\mathcal{X}$ of the instance $x^{*}$ which quantify how influential that feature was in the model predicting the value $h(x^{*})$.
@@
A widely used method for deriving attributions (importances) is the notion of SHapley Additive exPlanations, commonly referred to simply as SHAP values. Originally, the notion of Shapley values was introduced in the seminal work of Shapley [1952] in the context of cooperative game theory. The Shapley value is a mathematically well-founded and “fair” way of distributing a reward among all the members of a group playing a cooperative game and it is computed based on the rewards that would be received for all possible coalitions. The Shapley value is the unique way of distributing the reward that satisfies several reasonable mathematical properties that capture a notion of fairness [Shapley, 1952]. In the context of machine learning and statistics, the players become features, the reward is the prediction of the predictor $h$ and the SHAP value is the “contribution” or “influence” of that feature on the prediction. Shapley values are widely used due to their mathematical soundness and desirable properties Gromping [2007], Štrumbelj et al. [2009], Owen [2014], Datta et al. [2016], Owen and Prieur [2017], Lundberg and Lee [2017], Lundberg et al. [2020] Aas et al. [2021].
@@
Despite their prevalence, computing SHAP values is challenging, as it involves an exponential summation, i.e. a summation over exponentially many terms , see Equation 1. This is because the formula accounts for a feature’s importance in the context of all possible “coalitions” of other features, and therefore the formula covers all possible subsets of other features. Therefore, approximating them and speeding up the computation has received attention in a variety of settings. SHAP value computation can easily dominate the computation time of industry-level machine learning solutions on datasets with millions or more entries [Yang, 2021]. Yang [2021] point out that industrial applications sometimes require hundreds of millions of samples to be explained. Examples include feed ranking, ads targeting, and subscription propensity models. In these modeling pipelines, spending tens of hours in model interpretation becomes a significant bottleneck [Yang, 2021] and one usually needs to resort to multiple cores and parallel computing.
@@
Significant work has gone into speeding up the computation of SHAP values for a variety of settings. In the (ensemble of) trees setting, full access to the tree structure is assumed. Yang [2021], Bifet et al. [2022] provide theoretical and practical computational speedups to the classic TreeShap [Lundberg et al., 2020]. Similar to these results, in this work, we provide significant speedups for the tree setting over previous methods.
@@
As opposed to the tree setting, which is a “white box” setting, in the model-agnostic a.k.a black-box setting, we only have query access to the model. Here, our only means of access to the predictor is that we can pick an arbitrary $x\in\mathcal{X}^{n}$ and query the predictor for its value $h(x)$. The usual approach here is to approximate the exponential sum of the SHAP value computations using stochastic sampling [Covert and Lee, 2020, Mitchell et al., 2022a, Lundberg and Lee, 2017]. In this setting, Covert and Lee [2020], Mitchell et al. [2022a] provide sampling methods that require fewer queries to the black-box compared to vanilla KernelShap [Lundberg and Lee, 2017] for equal approximation accuracy. FastShap Jethani et al. [2021], which introduces a method for estimating Shapley values in a single forward pass using an end-to-end learned explainer neural network model. Our algorithm FourierShap falls into the query-access black-box setting as well.
@@
However, we take a different approach. We are guided by the key insight that many models used in practice have a “spectral bias”. Yang and Salman [2019], Valle-Perez et al. [2018] provably and experimentally show that fully connected neural networks with binary (zero-one) inputs learn low-degree – and therefore sparse – functions in a basis called the Walsh-Hadamard a.k.a Fourier basis. It is well known that the Walsh-Hadamard transform (WHT) of an ensemble of $T$ trees of depth $d$ is also of degree at most $d$ and moreover, $k=O(T4^{d})$-sparse [Kushilevitz and Mansour, 1993, Mansour, 1994].
@@
Our contributions: Guided by the aforementioned insights, we provide an algorithm to compute SHAP values, using the Fourier representation of the tree or black-box model. We first approximate the black-box function by taking its sparse Fourier transform. We theoretically justify, and show through extensive experiments, that for many real-world models such as fully connected neural networks and (ensembles of) trees this representation is accurate. Subsequently, we prove that SHAP values for a single Fourier basis function admit a closed-form expression not involving an exponential summation. Therefore, using the Fourier representation we overcome the exponential sum and can utilize compute power effectively to compute SHAP values. Furthermore, the closed-form expression we derive effectively “linearizes” the computation into a simple summation and is amenable to parallelization on multiple cores or a GPU. The Fourier approximation step is only done once, therefore FourierShap amortizes the cost of computing explanations for many inputs. Subsequently, SHAP computations using the Fourier approximation are orders of magnitude faster compared to other black-box SHAP approximation methods such as KernelShap and other variations of it which all involve a computationally expensive optimization. We show speedups over other methods such as DeepLift [Shrikumar et al., 2017] and FastShap [Jethani et al., 2021] as well. In addition to the speedup, our algorithm enables a reliable continuous trade-off between computation and accuracy, controlled in a fine manner by the sparsity of the Fourier approximation.
@@ 2.1 Shapley values
In game theory, a cooperative game is a function $v:2^{[n]}\rightarrow\mathbb{R}$ that maps a subset (coalition) $S\subseteq[n]$ of a group of players $[n]=\{1,\dots,n\}$ to their total reward (when they cooperate). If all the players cooperate they win a total reward of value $v([n])$. The main question is how they would distribute this reward among themselves. Shapley [1952] resolved this problem by a deriving a unique value based on “fairness axioms” proposed in his seminal work [Shapley, 1952]. The Shapley value of player $i\in[n]$ is:
@@
In the machine learning context, we have a predictor $h:\mathcal{X}^{n}\rightarrow\mathbb{R}$ mapping an $n$-dimensional feature vector to a value. In this context, the players become features $x_{i}\in\mathcal{X}$ and the reward is the prediction of the predictor $h$ and the Shapley value is the “contribution” or “influence” of the $i$’th feature on the prediction. We define $v(S)$ accordingly to capture this notion [Lundberg and Lee, 2017]:
@@
The choice of which distribution to average the missing features from has been discussed thoroughly in the relevant literature. As mentioned before, in this work, we focus on the SHAP values as defined in KernelShap introduced by Lundberg and Lee [2017], Lundberg et al. [2020] also known as “Interventional” [Janzing et al., 2020, Van den Broeck et al., 2021] or “Baseline” [Sundararajan and Najmi, 2020] SHAP, where the missing features are integrated out from the marginal distribution $p(\bm{x_{[n]\setminus S}})$, as opposed to the conditional distribution $p(\bm{x_{[n]\setminus S}}|\bm{x_{S}})$. We refer the reader to Appendix A.3 for a comprehensive overview of the literature discussing these two notions and their conceptual differences.
Since we will be using the well-known KernelShap [Lundberg and Lee, 2017] and its variant LinRegShap [Covert and Lee, 2020] as a baseline we briefly review their method here. Lundberg and Lee [2017] propose the “least squares characterization” of SHAP values. They prove that SHAP values are the solution to the following minimization problem:
@@
Unfortunately, the above optimization still involves an exponential sum. Therefore, Lundberg and Lee [2017] propose to sample subsets $S$ uniformly at random. Covert and Lee [2020], Mitchell et al. [2022a] provide better ways of sampling and solving the optimization to get approximations with lower variances and biases. Nevertheless, all these methods require solving a least squares minimization subject to constraints for each explained instance $x^{*}$ and, therefore, are computationally expensive.
Later, we also compare our method to FastSHAP [Jethani et al., 2021], a model-agnostic algorithm for computing SHAP values. In FastSHAP, a parametric explainer $\phi$ (e.g., MLP) is trained to directly generate SHAP values given data samples. Since training is only done once, this, similar to us, amortizes the cost of generating SHAP values across multiple instances. Computing SHAP values only requires a forward pass on the trained model and therefore can be done very quickly.
@@ 2.2 Fourier representations
Here we review the notions of the Fourier basis and what we mean by sparsity. We will later use sparse Fourier representation of functions to compute SHAP values. In this work, we focus on the setting where the inputs to the black-box function (predictor) are binary, i.e., $\mathcal{X}^{n}=\{0,1\}^{n}$. This means, we assume we have binary features and/or categorical features, through standard one-hot representations  [^1] Continuous features can be discretized into quantiles, enabling their transformation into categorical features.. The Fourier representation of the pseudo-boolean function $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ is the unique expansion of $h$ as follows: $h(x)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{f\in\{0,1\}^{n}}\widehat{h}(f)(-1)^{\langle f,x\rangle}$.
The inner product of two vectors $f,x\in\{0,1\}^{n}$ is defined as: $\langle f,x\rangle\equiv\sum_{i}^{n}f_{i}x_{i},\forall f,x\in\{0,1\}^{n}$.
@@
Assume $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ can be decomposed as follows: $h(x)=\sum\limits_{i=1}^{p}h_{i}(x_{S_{i}}),S_{i}\subseteq[n]$. That is, each function $h_{i}:\{0,1\}^{|S_{i}|}\rightarrow\mathbb{R}$ depends on at most $|S_{i}|$ variables. Then, $h$ is $k=O(\sum\limits_{i=1}^{p}2^{|S_{i}|})$-sparse and of degree $d=\max(|S_{1}|,\dots,|S_{p}|)$. (Proof in Appendix B.1)
@@
Let, $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be a pseudo-boolean function and let $d\in\mathbb{N}$ be some constant (w.r.t. $n$). If $h$ is of degree $d$, then, it is $k=O(n^{d})$-sparse. (Proof in Appendix B.1)
@@ 3.1 Spectral bias of fully connected neural networks
The function a fully connected neural network represents at initialization is a sample from a Gaussian Process (GP) [Rasmussen, 2004] in the infinite-width limit. Here, the randomness is over the initial weights and biases. The kernel $K$ of this GP is called the Conjugate Kernel (CK) [Daniely et al., 2016, Lee et al., 2017]. As mentioned before, here we investigate the case where the inputs to the neural network are binary, i.e., $\mathcal{X}^{d}=\{0,1\}^{d}$, similar to [Yang and Salman, 2019, Valle-Perez et al., 2018]. The CK kernel Gram matrix formed on the whole input space $\mathcal{X}^{d}=\{0,1\}^{d}$, has a simple eigenvalue decomposition: $\mathcal{K}\in\mathbb{R}^{2^{d}\times 2^{d}},\mathcal{K}=\sum\limits_{f\in\{0,1\}^{n}}\lambda_{f}u_{f}u_{f}^{\top},$ where $u_{f}\in\mathbb{R}^{2^{d}}$ is the eigenvector formed by evaluating the Fourier basis function $\Psi_{f}$ for different values of $x\in\{0,1\}^{n}$. Moreover, Yang and Salman [2019] show a weak spectral bias result in terms of the degree of $f$. Namely, the eigenvalues corresponding to higher degree frequencies have smaller values  [^2] To be more precise, they show that the eigenvalues corresponding to even and odd degree frequencies form decreasing sequences. That is, even and odd degrees are considered separately.. Given the kernel, $K$, and viewing a randomly initialized neural network function evaluated on the boolean $\{0,1\}^{n}$ as a sample from a GP one can see the following: This sample, roughly speaking, looks like a linear combination of the eigenvectors with the largest eigenvalues. This is due to the fact that a sample of the GP can be obtained as $\sum\limits_{i=1}^{2^{n}}\lambda_{i}\bm{w_{i}}u_{i},\bm{w_{i}}\sim\mathcal{N}(0,1)$. Combining this with the spectral bias results implies that neural networks are low-degree functions when randomly initialized.
@@
Going beyond neural networks at initialization, numerous studies have investigated the behavior of fully connected neural networks trained through (stochastic) gradient descent. Chizat et al. [2019], Jacot et al. [2018b], Du et al. [2018], Allen-Zhu et al. [2019a, b] found that the weights of infinite-width neural networks after training do not deviate significantly from their initialization, which has been dubbed "lazy training" by Chizat et al. [2019]. Lee et al. [2018, 2019] showed that training the last layer of an infinite-width randomly initialized neural network for an infinite amount of time corresponds to Gaussian process (GP) posterior inference with a certain kernel. Jacot et al. [2018b] extended the aforementioned results to training all the layers of a neural network (not just the final layer). They showed the evolution of an infinite-width neural network function can be described by the “Neural Tangent Kernel” [NTK, Jacot et al., 2018a] with the trained network yielding, on average, the posterior mean prediction of the corresponding GP after an infinite amount of training time. Lee et al. [2019] empirically showed the results carry over to the finite-width setting through extensive experiments. Yang and Salman [2019] again showed that the $u_{f}\in\mathbb{R}^{2^{d}}$ defined above are eigenvectors of the NTK Gram matrix and spectral bias holds. Gorji et al. [2023] validated these theoretical findings through extensive experiments in finite-width neural networks by showing that neural networks have “less tendency” to learn high-degree frequencies. We refer the reader to Appendix A.1 for a more comprehensive review.